# FuzzySQL: Uncovering Hidden Vulnerabilities in DBMS Special Features with LLM-Driven Fuzzing

### v1

@@ 3.4. Replay-Guided Crash Validation
To ensure that detected crashes are genuinely reproducible and rooted in state-dependent behaviors of the DBMS, FuzzySQL implements a replay-guided crash validation pipeline followed by test case reduction. This mechanism isolates the root cause of failures and extracts a minimal yet effective proof-of-concept (PoC) that reproduces the crash.
Replay-based crash validation. Unlike traditional fuzzers that fork a new process for each test case, FuzzySQL resets the database state using DROP DATABASE and CREATE DATABASE statements, preserving the global session context across executions. While this approach improves performance, it introduces the possibility of residual state effects. In many cases, a crash does not occur when the current test case is executed in isolation, but only when certain prior statements have altered the environment. To address this, FuzzySQL first attempts to re-execute the crashing test case alone. If the crash does not reproduce, it replays the full execution history, executing prior statements in order before the crashing input, to verify whether the failure is state-dependent. If the error only manifests during the full replay, the issue is classified as a state-dependent crash and proceeds to minimization.
@@
FuzzySQL’s replay and reduction workflow reconstructs the full crash path and eliminates irrelevant logic, producing a concise PoC of critical SQL statements, drastically improving debugging clarity and bug report quality.
@@ 5.1. Experimental Setup
To demonstrate the advantage of FuzzySQL, we compare it against three state-of-the-art open-source fuzzing baselines: Squirrel (Zhong et al., 2020), EET (Jiang and Su, 2024), and SQLancer (Rigger and Su, 2020a). Among them, Squirrel represents mutation-based fuzzing, while both EET and SQLancer adopt generation-based approaches. Since EET and SQLancer are not inherently a grey-box fuzzing tool, we first collect their generated test cases and then utilize FuzzySQL’s replay mechanism to measure its coverage. For SQLancer, we use TLP (Rigger and Su, 2020b) as the test oracle for MySQL, and NoREC (Rigger and Su, 2020a) for the other DBMS targets. This choice is constrained by the design of SQLancer, which employs different oracle-based testing strategies tailored to specific DBMSs. All baselines were evaluated under identical runtime conditions and DBMS versions. We adopted official configurations and standardized crash triage procedures to ensure fair and reproducible comparisons. Except for Squirrel, a mutation-based fuzzer using its default seeds, all baselines ran without a seed corpus. Each experiment was repeated five times on identical hardware, and the reported results are the averages, which show consistent statistical stability.
@@
MySQL 15 15 10 4
MariaDB 12 12 3 2
SQLite 2 2 1 1
ClickHouse 8 0 0 0
Total 37 29 14 7
@@ 5.2. Bug Discovery Effectiveness
To evaluate the practical effectiveness of FuzzySQL, we conducted a systematic fuzzing campaign on five representative open-source DBMSs. In total, FuzzySQL discovered 37 distinct bugs, among which 29 have been confirmed by vendors and 9 have been assigned CVE identifiers. As summarized in Table 2, 14 issues have already been fixed, while the remaining ones are scheduled for upcoming releases. Notably, 7 of the confirmed bugs are related to under-tested DBMS special features, such as GTID, PROCEDURE, and INSTALL COMPONENT, demonstrating FuzzySQL’s ability to explore feature-specific execution paths that are rarely exercised by existing fuzzers.
Beyond long-term bug discovery, we further compare fuzzing effectiveness under a fixed 24-hour budget. As shown in Figure 3, FuzzySQL consistently outperforms all baselines across all evaluated DBMSs in terms of newly discovered execution edges. The performance gap is particularly evident on MySQL, MariaDB, and PostgreSQL, where FuzzySQL achieves several times higher coverage than the strongest baseline. On SQLite and ClickHouse, the advantage is less pronounced, which can be attributed to the lightweight feature set of SQLite and the highly optimized execution engine of ClickHouse, where EET benefits from generating deeply nested analytical queries. Nevertheless, FuzzySQL still achieves the highest overall coverage across all systems.
To examine how this coverage advantage translates into concrete vulnerability discovery, Table 3 presents a breakdown of unique MySQL bugs discovered within 24 hours. Among the evaluated baselines, only EET uncovers a limited number of bugs, while none of the baselines discover any feature-related vulnerabilities. In contrast, FuzzySQL consistently exposes both feature-related and non-feature-related bugs within the same time budget. This result highlights a fundamental capability gap between FuzzySQL and existing fuzzers: while prior approaches can occasionally detect generic crashes, they lack the ability to synthesize and evolve stateful, feature-specific SQL programs required to exercise under-tested DBMS functionality.
Table 3. Unique MySQL bugs discovered by different fuzzers within 24 hours
@@
MySQL 13 $\bullet$ Assertion Failure
MySQL 15 $\circ$ NULL Ptr Deref
  • FR: $\bullet$ feature-related, $\circ$ not feature-related; !r: repair disabled; !m: mutation disabled.
  • $\dagger$ Bug ID corresponds to Table 4 (Appendix B).
Overall, these results demonstrate that FuzzySQL’s effectiveness lies not merely in discovering more bugs or achieving higher coverage, but in exposing qualitatively different classes of vulnerabilities that current DBMS fuzzers are structurally incapable of revealing. This capability stems from FuzzySQL’s integrated design, which enables efficient and systematic exploration of semantically sensitive logic across diverse database systems.
@@ 5.3. Ablation Study
Effectiveness of Logic Mutation. We analyze all 29 confirmed bugs discovered by FuzzySQL to identify those that depend on logic-shifting transformations. A bug is considered logic-sensitive if its PoC contains a condition (e.g., =, IN, JOIN) mutated from another semantically valid form and becomes non-triggering when reverted. Using this criterion, 20 out of 29 bugs (68.96%) are logic-sensitive, indicating that logic mutation is a major contributor to vulnerability discovery. Figure 5 summarizes the distribution of these bugs by transformation type. NULL-related rewrites account for the largest fraction (11 PoCs), followed by equality polarity changes (10 PoCs) and join rewrites (8 PoCs). Membership predicates (IN/NOT IN) and other transformations, including AND/OR rewrites, ordering reversals, and EXISTS manipulation, contribute additional unique triggers. Together, these results show that small, semantics-preserving logic shifts can effectively expose alternative execution paths that are rarely exercised by structure-preserving or generation-only fuzzers.
@@
Effectiveness of Automated Repair. Although semantic-aware repair introduces additional overhead, it substantially improves fuzzing yield. In our evaluation, 27% of the discovered bugs originate from test cases that were initially invalid and would have been discarded without repair, including cases with missing objects or unmet semantic preconditions.
Consistent with the system-level results and Table 3, disabling repair reduces the number of unique bugs discovered and causes several feature-related failures to be missed. These results indicate that automated repair is essential for preserving and exploiting high-value test cases, rather than serving as a mere convenience for handling invalid inputs.
@@ 5.4. Case Study
1CREATE TABLE t1 (v1 JSON, v2 CHAR(7), v3 BINARY(3));
2DESC ANALYZE FORMAT = JSON (SELECT v1 FROM t1 UNION XXXX);
@@
1KILL 1 IN (SELECT 1) IN (XXXX);
@@
1CREATE PROCEDURE test_proc (id VARCHAR(255)) BEGIN DECLARE dt
2DATETIME(6) DEFAULT ROW(XXXX) = SOME(SELECT 1) = ALL(XXXX);
@@ 6. Discussion
1INSTALL COMPONENT ’plugin2’ SET GLOBAL default.variable1 = ON, t0.c1 = 1 OR 1 = ANY (XXXX);
1CREATE TABLE t1 (XXXX);
2CREATE TABLE t2 (XXXX) WITHOUT ROWID;
3INSERT INTO t2 (XXXX) VALUES (XXXX);
4ANALYZE t2;
5WITH cte_0 AS (SELECT 1) SELECT DISTINCT ref_2.v1 FROM (cte_0 JOIN generate_series(79, 105, 7) AS ref_1 ON (ref_1.value = 1)) CROSS JOIN (t2 AS t2a JOIN (t1 AS ref_2 JOIN t2 ON (ref_2.v2 = 1))); -- Error: ambiguous column name: main.t2.v1
6WITH cte_0 AS (SELECT 1) SELECT DISTINCT ref_2.v1 FROM (cte_0 JOIN generate_series(79, 105, 7) AS ref_1 ON (ref_1.value = 1)) CROSS JOIN (t2 AS t2a JOIN (t1 AS ref_2 JOIN t2 as ref_3 ON (ref_2.v2 = 1))); -- Fixed and Trigger Crash
@@ 7.3. LLM-Assisted Fuzzing
8. Conclusion
We presented FuzzySQL, an LLM-driven fuzzing framework that uncovers deep-seated bugs in DBMS special features often missed by existing tools. By combining grammar-guided generation, semantic mutation, hybrid error repair, and state-aware execution, FuzzySQL enables broad feature coverage and precise bug reproduction. Evaluations across multiple DBMSs revealed 37 unique vulnerabilities, 7 of which were tied to under-tested DBMS features. These results validate the effectiveness of our approach and highlight the need for semantically guided fuzzing in database security. Our findings suggest that integrating LLMs with structured generation offers a promising direction for advancing semantic fuzzing across complex software systems.
Appendix
@@ Appendix A Prompts Used In FuzzySQL
{init_schema_statements}
@@
-- [Need to repair<
<SQL statement>
-- <error message>
-- (repair suggestion)
-- >Need to repair]
@@ Appendix B Vulnerability Status Across DBMSs
DBMS ID Reference Location Status
MySQL 1 CVE-redacted Optimizer Fixed
2 CVE-redacted Optimizer Fixed
3 CVE-redacted Optimizer Fixed
4 CVE-redacted Optimizer Fixed
5 CVE-redacted Optimizer Fixed
6 CVE-redacted Optimizer Fixed
7 CVE-redacted Optimizer Fixed
8 CVE-redacted Components Services Fixed
9 CVE-redacted Optimizer Fixed
10 S{redacted} Item_ref Fixed
11 S{redacted} Tmp_table Verified
12 Debug Build MDL_checker Verified
13 Debug Build Mysql_execute_command Verified
14 Debug Build Explain_query Verified
15 Debug Build Item_in_subselect Verified
SQLite 1 {redacted} WhereEnd Fixed
2 {redacted} VdbeExec Fixed
MariaDB 1 MDEV-redacted Optimizer Fixed
2 MDEV-redacted Tmp_table Fixed
3 MDEV-redacted Item_func_not Fixed
4 MDEV-redacted Item_subselect Verified
5 MDEV-redacted Item_subselect Verified
6 MDEV-redacted Item Verified
7 MDEV-redacted Item_field Verified
8 MDEV-redacted Filesort Verified
9 MDEV-redacted Select_value_catcher Verified
10 MDEV-redacted Window_funcs_sort Verified
11 MDEV-redacted Optimizer Verified
12 MDEV-redacted Filesort Verified
ClickHouse 1 {redacted} Condition_variable In progress
2 {redacted} FunctionIn In progress
3 {redacted} FunctionBinaryArithmetic In progress
4 {redacted} FormatStringHelperImpl In progress
5 {redacted} AlterCommand In progress
6 {redacted} ExecuteQueryImpl In progress
7 {redacted} Planner In progress
8 {redacted} IAST In progress
@@ 12 MDEV-redacted Filesort Verified
Status legend. We label an entry as Fixed once a patch is available in an upstream release; Verified indicates that we reproduced the issue and reported it, but an official fix is still pending; and In progress denotes that triage with the maintainers is ongoing. Overall, Table 4 summarizes 37 issues across four DBMSs (MySQL: 15, SQLite: 2, MariaDB: 12, ClickHouse: 8), with identifiers redacted where required by responsible disclosure.
Appendix C Open Science
This section follows the ACM CCS Open Science policy by listing the artifacts needed to evaluate our core contributions and explaining how reviewers can access them during double-blind review.
Access during double-blind review.
We provide an anonymized artifact repository at:
https://anonymous.4open.science/r/fuzzysql-D589/
Artifacts.
The repository contains the artifacts needed to evaluate this paper, including: (i) the implementation of FuzzySQL, (ii) experiment scripts/configurations, (iii) instructions to run the experiments.
Unavailable artifacts and responsible disclosure.
This work studies real-world security vulnerabilities. Detailed information about specific vulnerabilities (e.g., vulnerability-triggering inputs, proof-of-concept exploits, and affected systems) is not publicly disclosed at submission time due to responsible disclosure and security concerns. Several of the discovered vulnerabilities have been reported to the corresponding vendors and are undergoing coordinated disclosure. Releasing full vulnerability details prior to patch availability could pose security risks.
To enable evaluation of our methodology, we provide the full implementation of FuzzySQL, experiment scripts, configurations, and partially sanitized examples that illustrate the vulnerability patterns discussed in the paper, without exposing sensitive or actionable details.

### v2

@@ 3.4. Replay-Guided Crash Validation
To ensure that detected crashes are genuinely reproducible and to determine whether they depend on persistent DBMS state accumulated across executions, FuzzySQL implements a replay-guided crash validation pipeline followed by test case reduction. The goal of this stage is not only to confirm the existence of a crash, but also to identify whether the failure emerges from cross-statement or cross-execution state interactions and to extract a minimal yet effective proof-of-concept (PoC).
Replay-based crash validation. FuzzySQL deliberately preserves session-level execution history during fuzzing because many DBMS failures, especially those involving under-tested features, are not localized to a single standalone test case. Instead, they often emerge only after prior statements have modified internal metadata, runtime modes, replication state, or other persistent execution context. To support this behavior, FuzzySQL resets the logical database namespace using DROP DATABASE and CREATE DATABASE, while intentionally retaining the broader session context across executions. In this sense, replay-guided validation is not merely a debugging aid, but part of the bug-finding model of FuzzySQL.
When a crash is detected, FuzzySQL first re-executes the crashing test case in isolation to determine whether the failure is self-contained. If the crash does not reproduce, FuzzySQL replays the recorded execution history by re-executing prior statements in their original order before the crashing input. If the failure only manifests under such replay, the bug is classified as state-dependent. Rather than passively relying on incidental residual effects, FuzzySQL actively records crashing inputs together with their preceding execution history and validates them under controlled replay prefixes, so that failures rooted in persistent state interactions can be distinguished from purely local crashes.
@@
FuzzySQL’s replay and reduction workflow reconstructs the full crash path and eliminates irrelevant logic, producing a concise PoC of critical SQL statements, thereby improving debugging clarity and the quality of bug reports.
@@ 5.1. Experimental Setup
To demonstrate the advantage of FuzzySQL, we compare it against three state-of-the-art open-source fuzzing baselines: Squirrel (Zhong et al., 2020), EET (Jiang and Su, 2024), and SQLancer (Rigger and Su, 2020a). Among them, Squirrel represents mutation-based fuzzing, while both EET and SQLancer adopt generation-based approaches. Since EET and SQLancer are not inherently a grey-box fuzzing tool, we first collect their generated test cases and then utilize FuzzySQL’s replay mechanism to measure its coverage. For SQLancer, we use TLP (Rigger and Su, 2020b) as the test oracle for MySQL, and NoREC (Rigger and Su, 2020a) for the other DBMS targets. This choice is constrained by the design of SQLancer, which employs different oracle-based testing strategies tailored to specific DBMSs. All baselines were evaluated under identical runtime conditions and DBMS versions. We adopted official configurations and standardized crash triage procedures to ensure fair and reproducible comparisons. Except for Squirrel, a mutation-based fuzzer using its default seeds, all baselines ran without a seed corpus. Each experiment was repeated 5 times on identical hardware, and the reported results are the averages, which show consistent statistical stability.
@@
MySQL 22 22 12 10
MariaDB 28 24 10 9
SQLite 2 2 2 1
ClickHouse 12 12 7 7
Total 64 60 31 27
@@ 5.2. Bug Discovery Effectiveness
To evaluate the practical effectiveness of FuzzySQL, we conducted a systematic fuzzing campaign on five representative open-source DBMSs. In total, FuzzySQL discovered 64 distinct bugs, among which 60 have been confirmed by vendors and 9 have been assigned CVE identifiers. As summarized in Table 2, 31 issues have already been fixed, while the remaining ones are scheduled for upcoming releases. Notably, 27 of the confirmed bugs are related to under-tested DBMS special features. This observation suggests that FuzzySQL does not merely improve generic SQL fuzzing coverage, but is particularly effective at exposing failures rooted in feature-specific and under-explored DBMS behaviors.
Figure 3. The number of new edges discovered in 24 hours.
Beyond long-term bug discovery, we further compare fuzzing effectiveness under a fixed 24-hour budget. As shown in Figure 3, FuzzySQL consistently outperforms all baselines across the evaluated DBMSs in terms of newly discovered execution edges. The advantage is especially pronounced on MySQL, MariaDB, and PostgreSQL, where FuzzySQL achieves substantially higher coverage than the strongest baseline. On SQLite and ClickHouse, the margin is smaller. This is likely because SQLite exposes a relatively lightweight feature space, while ClickHouse already benefits from EET’s generation of deeply nested analytical queries. Even so, FuzzySQL still attains the highest overall coverage across all five systems.
To understand how this coverage advantage translates into concrete vulnerability discovery, Table 3 reports deduplicated MySQL bugs discovered within 24 hours and shows which fuzzers were able to trigger each bug. Among the evaluated baselines, only EET uncovers a limited number of bugs, while none of the baselines expose feature-related vulnerabilities in this setting. In contrast, FuzzySQL consistently triggers both feature-related and non-feature-related bugs under the same time budget. This result indicates an important capability gap between FuzzySQL and prior DBMS fuzzers: although existing approaches can occasionally detect generic crashes, they are considerably less effective at constructing and evolving the semantically valid, state-sensitive SQL sequences required to exercise under-tested DBMS functionality.
Overall, these results show that the effectiveness of FuzzySQL lies not only in finding more bugs or reaching higher coverage, but also in exposing a broader and qualitatively different set of vulnerabilities. In particular, feature-related bugs are not isolated anecdotes in our results; with 27 confirmed cases, they constitute a substantial portion of the bugs uncovered by FuzzySQL. This capability arises from FuzzySQL’s integrated design, which jointly addresses three obstacles in feature-related DBMS bug discovery: insufficient coverage of feature-specific constructs, unmet semantic preconditions, and hidden state interactions that span multiple statements or executions.
Table 3. Deduplicated MySQL bugs discovered within 24 hours
@@
MySQL 17 $\bullet$ Assertion Failure
MySQL 21 $\circ$ NULL Ptr Deref
  • $\dagger$ Bug ID corresponds to Table 4 (Appendix B). FR: $\bullet$ feature-related, $\circ$ not feature-related; !r: repair disabled; !m: mutation disabled.
@@ 5.3. Ablation Study
Effectiveness of Logic Mutation. We analyze all 60 confirmed bugs discovered by FuzzySQL to identify those that depend on logic-shifting transformations. A bug is considered logic-sensitive if its PoC contains a condition (e.g., =, IN, JOIN) mutated from another semantically valid form and becomes non-triggering when reverted. Using this criterion, 22 out of 60 bugs (36.67%) are logic-sensitive, indicating that logic mutation is a major contributor to vulnerability discovery. Figure 5 summarizes the distribution of these bugs by transformation type. Equality polarity changes account for the largest fraction (10 PoCs), followed by JOIN-related rewrites (9 PoCs) and membership predicates IN-related rewrites (6 PoCs). Other transformations, including AND/OR rewrites, ordering reversals, and EXISTS manipulation, contribute additional unique triggers. Together, these results show that small, semantics-preserving logic shifts can effectively expose alternative execution paths that are rarely exercised by structure-preserving or generation-only fuzzers.
@@
Effectiveness of Automated Repair. Although semantic-aware repair introduces additional overhead, it substantially improves fuzzing yield. In our evaluation, 16 out of the discovered bugs originate from test cases that were initially invalid and would have been discarded without repair, including cases with missing objects or unmet semantic preconditions. Consistent with the system-level results and Table 3, disabling repair reduces the number of unique bugs discovered and causes several feature-related failures to be missed. These results indicate that automated repair is essential for preserving and exploiting high-value test cases, rather than serving as a mere convenience for handling invalid inputs.
@@ 5.4. Case Study
1 CREATE TABLE t1 (v1 JSON, v2 CHAR(7), v3 BINARY(3));
2 DESC ANALYZE FORMAT = JSON (SELECT v1 FROM t1 UNION XXXX);
@@
1 KILL 1 IN (SELECT 1) IN (XXXX);
@@
1 CREATE PROCEDURE test_proc (id VARCHAR(255)) BEGIN DECLARE dt
2 DATETIME(6) DEFAULT ROW(XXXX) = SOME(SELECT 1) = ALL(XXXX);
@@
1 INSTALL COMPONENT ’plugin2’ SET GLOBAL default.variable1 = ON, t0.c1 = 1 OR 1 = ANY (XXXX);
1 CREATE TABLE t1 (XXXX);
2 CREATE TABLE t2 (XXXX) WITHOUT ROWID;
3 INSERT INTO t2 (XXXX) VALUES (XXXX);
4 ANALYZE t2;
5 WITH cte_0 AS (SELECT 1) SELECT DISTINCT ref_2.v1 FROM (cte_0 JOIN generate_series(79, 105, 7) AS ref_1 ON (ref_1.value = 1)) CROSS JOIN (t2 AS t2a JOIN (t1 AS ref_2 JOIN t2 ON (ref_2.v2 = 1))); -- Error: ambiguous column name: main.t2.v1
6 WITH cte_0 AS (SELECT 1) SELECT DISTINCT ref_2.v1 FROM (cte_0 JOIN generate_series(79, 105, 7) AS ref_1 ON (ref_1.value = 1)) CROSS JOIN (t2 AS t2a JOIN (t1 AS ref_2 JOIN t2 as ref_3 ON (ref_2.v2 = 1))); -- Fixed and Trigger Crash
@@ Appendix A Prompts Used In FuzzySQL
ΨΨ{init_schema_statements}
Ψ
@@
$Ψ$-- [Need to repair<
ΨΨ<SQL statement>
$Ψ$-- <error message>
$Ψ$-- (repair suggestion)
$Ψ$-- >Need to repair]
Ψ
@@ Appendix B Vulnerability Status Across DBMSs
DBMS ID Reference Location Feature-related Status
MySQL 1 CVE-redacted Optimizer $\circ$ Fixed
2 CVE-redacted Optimizer $\circ$ Fixed
3 CVE-redacted Optimizer $\circ$ Fixed
4 CVE-redacted Optimizer $\circ$ Fixed
5 CVE-redacted Optimizer $\circ$ Fixed
6 CVE-redacted Optimizer $\circ$ Fixed
7 CVE-redacted Optimizer $\circ$ Fixed
8 CVE-redacted Components Services $\bullet$ Fixed
9 CVE-redacted Optimizer $\circ$ Fixed
10 S{redacted} Item_ref $\circ$ Fixed
11 S{redacted} Tmp_table $\circ$ Verified
12 S{redacted} Sql_executor $\bullet$ Fixed
13 S{redacted} Table_ref $\circ$ Fixed
14 S{redacted} Item_tree $\bullet$ Verified
15 S{redacted} Mutex $\bullet$ Verified
16 S{redacted} Log_message $\bullet$ Verified
17 Debug Build Mysql_execute_command $\bullet$ Verified
18 Debug Build MDL_checker $\bullet$ Verified
19 Debug Build Explain_query $\bullet$ Verified
20 Debug Build Sp_head $\bullet$ Verified
21 Debug Build Item_in_subselect $\circ$ Verified
22 Debug Build User_table $\bullet$ Verified
MariaDB 1 MDEV-redacted server $\bullet$ Verified
2 MDEV-redacted Optimizer $\circ$ Fixed
3 MDEV-redacted Optimizer $\circ$ Fixed
4 MDEV-redacted Temporal Types $\circ$ Verified
5 MDEV-redacted Server $\bullet$ Fixed
6 MDEV-redacted Server $\bullet$ Verified
7 MDEV-redacted Server $\circ$ Fixed
8 MDEV-redacted Optimizer $\circ$ Verified
9 MDEV-redacted Optimizer $\circ$ Fixed
10 MDEV-redacted Optimizer $\circ$ Verified
11 MDEV-redacted Optimizer $\circ$ Verified
12 MDEV-redacted Stored routines $\bullet$ Verified
13 MDEV-redacted Server $\bullet$ Verified
14 MDEV-redacted Server $\bullet$ Fixed
15 MDEV-redacted server $\bullet$ Fixed
16 MDEV-redacted Optimizer $\circ$ Verified
17 MDEV-redacted Optimizer $\circ$ Verified
18 MDEV-redacted Optimizer $\circ$ Verified
19 MDEV-redacted Server $\bullet$ Fixed
20 MDEV-redacted Optimizer $\circ$ Verified
21 MDEV-redacted Server $\circ$ Fixed
22 MDEV-redacted Data types $\circ$ Verified
23 MDEV-redacted Stored routines $\bullet$ Fixed
24 MDEV-redacted Optimizer $\circ$ Verified
25 MDEV-redacted Add_key_field $\circ$ In progerss
26 MDEV-redacted Item_direct_view_ref $\circ$ In progerss
27 MDEV-redacted Item_func_or_sum $\circ$ In progerss
28 MDEV-redacted Setup_copy_fields $\circ$ In progerss
SQLite 1 {redacted} WhereEnd $\bullet$ Fixed
2 {redacted} VdbeExec $\circ$ Fixed
ClickHouse 1 {redacted} FunctionIn $\circ$ Verified
2 {redacted} FunctionBinaryArithmetic $\circ$ Fixed
3 {redacted} InterpreterCreateQuery $\circ$ Verified
4 {redacted} Modify $\bullet$ Fixed
5 {redacted} ExecuteQueryImpl $\bullet$ Verified
6 {redacted} Planner $\bullet$ Fixed
7 {redacted} GetColumnName $\circ$ Fixed
8 {redacted} setAlias $\bullet$ Verified
9 {redacted} IdentifierResolveScope $\bullet$ Fixed
10 {redacted} ExpressionActions $\circ$ Fixed
11 {redacted} Rename $\bullet$ Verified
12 {redacted} AssertTypeEquality $\bullet$ Fixed
@@ 28 MDEV-redacted Setup_copy_fields $\circ$ In progerss
Status legend. We label an entry as Fixed once a patch is available in an upstream release; Verified indicates that we reproduced the issue and reported it, but an official fix is still pending; and In progress denotes that triage with the maintainers is ongoing. Overall, Table 4 summarizes 64 issues across four DBMSs (MySQL: 22, MariaDB: 28, SQLite: 2, ClickHouse: 12), with identifiers redacted where required by responsible disclosure.